OpenStudy (anonymous):

what is an equation for the line that passer through points (-1,-4 and (1,4) write the equation is slope intercept form

OpenStudy (anonymous):

you almost done writing ?

OpenStudy (texaschic101):

we need to find the slope first by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-1,-4) x1 = -1 and y1 = -4 (1,4) x2 = 1 and y2 = 4 now we sub slope(m) = (4 - (-4)) / (1 -(-1)) m = (4 + 4) / (1 + 1) m = 8/2 m = 4 now that we know the slope is 4 and we have a set of points, we will use the point slope formula: y - y1 = m(x - x1) slope(m) = 4 you can use either of your sets of points...I will use (1,4) (1,4) x1 = 1 and y1 = 4 now lets sub y - 4 = 4(x - 1) -- distribute through the parenthesis y - 4 = 4x - 4 -- add 4 to both sides y = 4x - 4 + 4 y = 4x any questions ?

OpenStudy (anonymous):

no i think i have it that you

OpenStudy (texaschic101):

no problem :)